OpenStudy (anonymous):

Can anyone shed some light on evaluating a piecewise-defined function below it is asking for 3 answers? Question: Suppose that the function h is defined, for all real numbers, as follows: all of the number are enclosed on one side in one big bracket I can’t draw it. h(x) = {-1/2 x -2 if x ≤ -2 {(x-1)^2 -2 if -2 < x ≤ 1 {-2 if x > 1 Find the following: h(-3) = ___ h(-2) = ___ h(-1) = ___

OpenStudy (anonymous):

It just wants you to draw: the line -1/2 x -2 for values less than and equal to -2 (x-1)^2 -2 for x values between -2 and 1 and the line y=2 for x greater than 1 =) h(-3) ..now which region does -3 come in ? -3 <-2? Yes true! So the value h(-3) we need is -1/2 (-3) -2 = 3/2 -2 = -1/2 Hope this helps you get started!

OpenStudy (anonymous):

Thank you! Here is your medal!

OpenStudy (anonymous):

You are soooo welcome! Glad I was helpful <3